OpenStudy (anonymous):

Omg i think the anwser is 9 but what is 6/2(1+2)=?

OpenStudy (anonymous):

1 because 6/6 is 1.

OpenStudy (anonymous):

6/2(1+2)=0 6/2(3)=0 6/6=0 The answer is 1

OpenStudy (anonymous):

Dont YOU USe P.E.M.D.A.S i thought u do 1+2 then 6/2 then times 3 and 3 wich equals nine

OpenStudy (anonymous):

If you are using PEMDAS it would be (1+2)=3 them multiplication so 3x2= 6 then division which is 6/6=1

OpenStudy (anonymous):

PEMDAS 6/2(1+2) 6/2(3) 6/6 1 you multiply before you divide

OpenStudy (anonymous):

left to right

OpenStudy (anonymous):

the answer is 1 

OpenStudy (zarkon):

9

OpenStudy (zarkon):

\[6/2(1+2)=6\times\frac{1}{2}\times(1+2)\]

OpenStudy (goldphenoix):

It's 9, wrongy. Everyone above me is wrong, except Zarkon.

OpenStudy (anonymous):

No because the "/" means division.

OpenStudy (goldphenoix):

So how would you do it?

OpenStudy (anonymous):

its the equivalent of saying \[6\div(2(1+2))\]

OpenStudy (goldphenoix):

6/2(1+2) can be written as: \[\large \large \frac{ 6 }{ 2 } (1+2)\] Correct so far?

OpenStudy (anonymous):

No. Following PEMDAS, you would first add 1 and 2 together to get 3. Then you would multiply 3 by 2 and get six. finally, you would divide 6 by 6 and get 1. Make sense?

OpenStudy (goldphenoix):

The first step is right. You must do parenthesis first. So 1 + 2 is 3. Next, you have division or multiplication, right?

OpenStudy (anonymous):

you do multiplication first. you can remember that by P-E-Multiplication-division-A-S

OpenStudy (goldphenoix):

I guess you don't know what PEMDAS means.

OpenStudy (goldphenoix):

PEMDAS = Parenthesis Exponent Multiplication/Division Division/Multiplication Addition/Subtraction Subtraction/Addition.

OpenStudy (zarkon):

multiplication and division is done from left to right

OpenStudy (goldphenoix):

What Zarkon said.

OpenStudy (goldphenoix):

http://www.mathsisfun.com/operation-order-pemdas.html <-- Learn more about PEMDAS.

OpenStudy (anonymous):

If division was first then is would be spelled PEDMAS

OpenStudy (zarkon):

PEMDAS is just a way to help you...it does not give the full picture

OpenStudy (goldphenoix):

...

OpenStudy (goldphenoix):

What Zarkon said. Anyway, google PEMDAS. Everything will say Multiplication or division. (Whichever comes first from left to right)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6%2F%282%281%2B2%29%29 Here wolfram knows the answer

OpenStudy (zarkon):

you didnt type it correctly.

OpenStudy (zarkon):

type it as is

OpenStudy (zarkon):

you will get 9

OpenStudy (goldphenoix):

The input is wrong.

OpenStudy (anonymous):

So it could go either way depending on where the parenthesis are?

OpenStudy (goldphenoix):

\[\large \large \frac{ 6 }{ 2(1+2) } \neq \frac{ 6 }{ 2 } (1+2)\]

OpenStudy (zarkon):

\[6/2(1+2)=6\div 2(1+2)=6\frac{1}{2}(1+2)\] it will always be 9

OpenStudy (goldphenoix):

Another way of writing 6/2 (1+2) is: \[\large \large \frac{ 6 (1+2)}{ 2 } \ \] Does it make sense?

OpenStudy (anonymous):

Ugh I feel bad that I gave her that answer then :(

OpenStudy (zarkon):

she will live :)

OpenStudy (anonymous):

@piercethesyd ik it was 9 i just wanted to make sure i was rite i got it rite but thnx